OpenStudy (anonymous):

Four times a certain number is the same as the number increased by 78. Find the number

OpenStudy (anonymous):

dude that's to easy

OpenStudy (anonymous):

Let's say the number is x. Then 4x=x+78. Then you solve for x.

OpenStudy (anonymous):

subtract 1x into 4x

OpenStudy (anonymous):

4(number) = (number) + 78

OpenStudy (anonymous):

dude r u going to reply

OpenStudy (anonymous):

What if i get decimel?

OpenStudy (anonymous):

The result is an integer.

OpenStudy (anonymous):

i got an decimal to ???

OpenStudy (anonymous):

i got 19.5

OpenStudy (anonymous):

i divided 3 into 78

OpenStudy (anonymous):

and i got 0.038

OpenStudy (anonymous):

\[4x=x+78\]\[4x-x=78\]\[3x=78\]\[x=\frac{78}{3}\] (@dreap3 You said that the problem was too easy. Though you don't know how to solve it.)

OpenStudy (anonymous):

@trev2673 Now do the division and you will get the result. Please tell me the result when you get it.

OpenStudy (anonymous):

26

OpenStudy (anonymous):

Good job trev! Did you understand the procedure?

OpenStudy (anonymous):

dude i know how to do it i set the formula wrong............. i did 3 /78

OpenStudy (anonymous):

Yes it really did.Thank-you, I appreciate it.